OpenStudy (booklover1997):

Binomial Theorem help! Show the work to find the third term of (2x-3y)^7. Then use your calculator to help find the term.

zepdrix (zepdrix):

Lol I've seen this exact question posted several times in the past few days :) I wonder if you're all in the same class or something hehe.

zepdrix (zepdrix):

Ok so Binomial Theorem:

zepdrix (zepdrix):

\[\large\rm (\color{orangered}{a}+\color{royalblue}{b})^n\quad=\sum_{k=0}^n \left(\begin{matrix}n \\ \rm k\end{matrix}\right)\left(\color{orangered}{a}\right)^{n-k}\left(\color{royalblue}{b}\right)^k\]

zepdrix (zepdrix):

So for our specific problem we have:\[\large\rm (\color{orangered}{2x}+\color{royalblue}{-3y})^7\quad=\sum_{k=0}^7 \left(\begin{matrix}7 \\ \rm k\end{matrix}\right)\left(\color{orangered}{2x}\right)^{7-k}\left(\color{royalblue}{-3y}\right)^k\]

zepdrix (zepdrix):

If you ignore the summation, then this gives us the expression for our `general term`.\[\large\rm general~term:\qquad \left(\begin{matrix}7 \\ \rm k\end{matrix}\right)(2x)^{7-k}(-3y)^k\]

zepdrix (zepdrix):

Recall that k starts counting from 0, not from 1! So our first term is the 0th term, our second term being the k=1 term, and our third term being the k=2 term. So our third term will correspond to k=2, not k=3!

zepdrix (zepdrix):

So plug k=2 into your general term, and simplify.